# Multi-Source EEG Emotion Recognition via Dynamic Contrastive Domain Adaptation

### v1

@@ V-A Cross-Subject Results
The comparisons of the emotion recognition results of our MS-DCDA model and the baseline algorithms are presented in Tables I, where the best results are highlighted in bold font. The alternative methods are tuned to their best performances. If the relevant parameters are published, we will use them. If not, the parameters will remain consistent with our method. The cross-subject experimental results on SEED and SEED-IV, which are based on both the state-of-the-art methods and our proposed method, are presented in Table II. On the SEED dataset, the constructed MS-DCDA model significantly improves recognition performance, achieving 90.84% for the third session and 88.75% for the average of three sessions. In comparison to other methods, such as MS-MDA and MS-ADA, both being multi-source domain adaptation methods, our method demonstrates superior performance. Compared to MS-MDA which achieves the highest average accuracy, our method has improved accuracy by 3.01%. Compared to MS-ADA which achieves the highest average accuracy, our method has improved accuracy by 3.77%. On the SEED-IV dataset, our method achieved the highest accuracy 78.49% for the second session, which is 2.38% higher than MS-MDA, 74.34% for the average of three sessions, which is 3.12% higher than MS-MDA. Additionally, both F1 and Kappa scores of our model outperform the alternative methods.
@@
To further validate the superiority of our method, we compared it with another set of state-of-the-art methods. As illustrated in Table III, our method outperforms other methods in mean accuracy for both the SEED and SEED-IV datasets. The multi-source domain adaptation algorithm MFA-LR, HVF2N-DBR, and MWACN also work well in the cross-subject experiments. Our algorithm dynamically balance the model domain transferability and discriminability and achieved higher accuracy than MFA-LR (uses class-level alignment loss) and MWACN (uses association reinforcement to adapt conditional distribution). The recognition accuracy is respectively increased by 1.51% and 3.50% in cross-subject experiments compared to the alternative methods (HVF2N-DBR for SEED dataset and MFA-LR for SEED-IV dataset).
TABLE VI: ABLATION STUDY ON LOSS FUNCTIONS
@@
CE only 84.13±7.89 90.68±6.90
w/o. MMD 88.01±9.95 93.41±5.87
w/o. DISC 86.65±9.61 92.16±5.63
w/o. SCD 84.26±7.66 90.95±5.00
ALL 88.75±8.88 93.92±8.01
TABLE VII: ABLATION STUDY ON STATIC vs. DYNAMIC
LOSS RATIO $L_{DISC}$ : $L_{MMD}$
@@
TABLE VIII: ABLATION STUDY ON NET LAYER NORMALIZATION (LN)
Method Cross-subject Cross-session SEED SEED-IV SEED SEED-IV CFE+LN 85.51±7.99 70.71±9.45 93.59±7.52 76.12±14.16 MBC+LN 82.86±7.49 70.89±11.07 90.06±7.21 77.46±12.32 BOTH+LN 88.75±8.89 74.34±10.62 93.92±8.01 77.53±12.13
TABLE IX: ABLATION STUDY ON BRAIN LOBES
Lobes Cross-subject Cross-session SEED SEED-IV SEED SEED-IV F 86.56±9.63 73.5±14.09 92.97±7.64 75.57±12.50 O 85.14±8.76 66.03±9.73 91.27±7.93 73.64±12.53 P 85.62±9.17 70.41±11.60 92.35±7.37 76.94±12.62 T 77.93±13.19 65.45±11.81 87.78±9.37 71.31±13.59 F+P 87.53±9.57 74.12±11.20 95.22±5.71 76.94±12.86 O+T 85.55±10.84 72.10±13.22 91.91±8.02 74.86±12.07 F+P+O 86.35±8.85 72.25±12.49 92.11±7.50 77.64±13.10 F+P+T 88.33±8.78 72.55±13.26 94.57±6.10 78.15±12.02 All 88.75±8.89 74.34±10.62 93.92±8.01 77.53±12.13 F: Frontal; O: Occipital; P: Parietal; T: Temporal
V-B Cross-Session Results
The specific subject results of the best session in cross-subject and cross-session experiments are shown in Tables IV. It can be seen that in various experiments, our algorithm achieved the highest accuracy with the highest number of subjects compared to other algorithms. Furthermore, besides our method, DAN, MS-MDA, and MS-ADA all achieved the best results on certain subjects in both cross-subject and cross-session experiments. Our method integrates dynamic, multi-source domain adaptation, and class alignment concepts, amalgamating the strengths of these three methods and yielding superior performance. On the SEED dataset, some subjects achieved the highest accuracy of 100% in cross-subject and cross-session experiments. On the SEED-IV dataset, the subject achieved the highest accuracy of 95.55% and 94.98% in cross-subject and cross-session experiments. Notably, the results on the SEED dataset are significantly better than the results on the SEED-IV dataset. On the one hand, this may be because the SEED-IV dataset has more sentiment classification than the SEED dataset, which increases the difficulty of class feature recognition. On the other hand, it may be because the sample size of SEED-IV is much smaller than that of SEED. The cross-session experimental results of SEED and SEED-IV are provide in Table V. Our MS-DCDA model outperformed the rest models for all four metrics.
Figure 3: Confusion matrices analysis of our MS-DCDA model for (a) Cross-subject experiment on SEED dataset; (b) Cross-subject experiment on SEED-IV dataset; (c) Cross-session experiment on SEED; and (d) Cross-session experiment on SEED-IV. Our model shows the strongest sensitivity to positive and neutral emotions respectively for the SEED and SEED-IV dataset, with slightly reduced sensitivity for the rest emotions.
V-C Ablation Studies
In this section, we conduct ablation experiments to investigate the different components and equilibrium factors of the model on performance. For the sake of conciseness and consistency, all ablation experiments were conducted on the SEED. The results of the ablation experiment for the model components in cross-subject and cross-session experiments are shown in Table VI. Compared to the accuracy of 88.75% for full model components in cross-subject experiments, the removal of the MMD loss leads to a decrease in accuracy by 0.74%, omitting the disc loss decreases accuracy by 2.1%, and excluding the SCD loss results in a decrease of 4.49%. It can be seen that SCD has the greatest performance improvement, while MMD has the smallest performance improvement. Even without using SCD, the accuracy remains higher than the baseline model by 0.13%. Compared to the accuracy of 93.92% for full model components in cross-session experiments, excluding the MMD loss results in a decreased accuracy of 0.51%, removing the disc loss reduces accuracy by 1.76%, and omitting the SCD loss leads to a decrease of 2.97%. Similar to the cross-subject experiments, SCD shows the most substantial performance improvement, followed by disc, and MMD. The accuracy of not using SCD is still higher than that of the baseline model by 0.27%. These results emphasize the indispensability of each loss component. Table VII presents the comparison of using dynamic weighting factor $\tau$ in Eq. 14 against several combinations of static weights. It can be seen that using dynamic weighting factor shows better performance. Using dynamic weighting factor $\tau$ promoting dynamic balance between domain alignment and class discrimination effectively avoids performance degradation caused by excessive pursuit of alignment or discriminability.
V-D Visual Analysis
Confusion Matrix Analysis. The confusion matrices of cross-subject and cross-session experiments for three sessions on SEED and SEED-IV are provide in Fig. 3. As can be seen in Fig. 3a and 3b, our MS-DCDA method exhibits strongest sensitive to the positive emotion on the SEED data, with slightly reduced sensitivity in neutral and negative emtions. It is observed in Fig. 3c and 3d that neutral emotion is less likely being confused than fear and sadness on SEED-IV by our model, which may be explain by the fact that both fear and sadness are particularly negative and exhibit similarities in the activated electrode signals.
T-SNE Analysis. The t-sne plots [59] of feature distributions are provide in Fig. 4 for SEED dataset, which is richer in source diversity. We randomly picked 100 EEG samples from each subject for visualization to display changes in feature distribution caused by model training. As shown in Fig. 4a, the distribution of original features, it can be seen that most of the samples are concentrated in a singular area, accompanied by a few outliers among individual subjects. The observed feature distribution validates our hypothesis that all EEG data possess certain low-level features, with their distributions in the feature space exhibiting slight overlap. Following normalization, this phenomenon becomes more pronounced, resulting in a more uniform and centralized distribution of EEG data. Fig. 4b and Fig. 4c show the feature distributions processed by the DAN model and MS-DCDA model, respectively. DAN brings the source domain features and target domain features closer together, creating an overlap. On the other hand, MS-DCDA represents each source domain and target domain feature distribution separately, resulting in a certain degree of similarity and overlap between the targets and each individual source domain. Fig. 4d shows the initial distribution of combined source domain features and target domain features. We also show the feature distributino of source ensemble and target learned by DAN and our MS-DCDA respectively. While maintaining the consistency in source domain, it can be observed in Fig. 4e that DAN does not show a clear trend of class clutering. In contrast, our MS-DCDA model adapts T to individual sources and a complementary multi source ensemble (SE) with class-awareness. It also dynamically adjusts the weights of domain transferability and discriminability, leading to improved classification accuracies, wider inter-class margin, and higher intra-class compactness (see Fig. 4f).
Data Transfer Studies. The SEED and SEED-IV datasets are similar in experimental design, collection methods, data pre-processing, and analogical categories. They differ only in the number of emotion classes. Here we train our model with one dataset and test using the other as dataset transfer studies. To have the same number of emotion classes in each set, we merge the sad and fear classes in SEED-IV into a single negative class in SEED, and correspond neutral and happy in SEED-IV with the neutral and positive in SEED respectively. The results of the data transfer studies are shown in Fig. 5. It is observed that our MS-DCDA model outperforms alternative SS-DA and MS-DA algorithms in both cross-subject (see Fig. 5a) and cross-session experiments (see Fig. 5b), demonstrating an improved generalization ability. It is also found that the effect of data order (e.g., forward vs. backward) on the performance of data transfer is minimal, especially in cross-session experiments, which may be explained by the similarity of the two datasets. The inter-class and intra-class alignment in our MS-DCDA model reveals the fine-grained emotional features, enhancing the robustness of our model in various transfer scenarios.
Figure 4: T-SNE illustration of domain adaptation and emotion recognition on SEED dataset. (a) Distribution of 14 individual source subjects (S1-S14), the ensemble of the 14 individual sources (S15, source ensemble) and a target subject. (b) Distribution of the single source (S) and the target learned by DAN [35]. (c) Distribution of the 14 individual source subjects and the target subject learned by our MS-DCDA. (d) The Distribution of the ensemble of the source ensemble and the target. (e) Distribution of the single source and the target learned by DAN [35]. (f) Distribution of the source ensemble and the target learned by our MS-DCDA.
VI DISCUSSION
We evaluate the proposed framework on the SEED and SEED-IV datasets, using the leave-one-out cross-validation for cross-subject and cross-session experiments. The MS-DCDA framework demonstrated improved emotion recognition performance on both datasets. In cross-subject experiments, we achieved an accuracy of 90.48% and 90.85% for sessions 1 and 3 on the SEED dataset, with a relatively lower accuracy of 84.98% in session 2. On the SEED-IV dataset, our MS-DCDA performs better in session 2 with a leading 78.49% accuracy, while the accuracies are slightly reduced to 72.36% and 72.16% for sessions 1 and 3. The variance in performance between sessions indicates significant data dependency where samples from other sessions may exhibit undesired transfer. The cross-session variations can result in variations in recognition accuracies if data are obtained through repetitive experiments. To avoid this issue, we conducted extensive cross validation to demonstrate the robustness of the proposed framework. The observation that cross-subject results appear to be less robust compared to the corresponding cross-session results may be explained by the individual subject disparities in EEG being more complicated than its non-stationary problems.
As demonstrated in in Tables II, III and V, multi-source domain adaptation generally achieve superiorier performance than single-source methods for EEG-based emotion recognition tasks. The ablation studies on various loss functions shown in Table VI, where the accuracy of our model improved from 84.13% and 90.68% by incorporating multi-source loss penalties reaffirms this observation. The use of one-to-one neural network branches for each pair of source and target allows adaptation of target to sources of different distributions through comprehensive feature representations, thereby improving the model’s generalization ability and performance. However, computational complexity and training time may increase linearly with the number of sources. Using a relatively shallow and simple neural network structure, such as a few fully connected layers, may help alleviate the training cost.
@@ VI DISCUSSION
The proposed MS-DCDA successfully maintain the coarse-grained inter-domain alignment while improving the intra-domain and intra-class alignment through the fine-grained class-level adapation. With the use of the SCD penalty, the increases in recognition accuracy of 4.49% and 2.97% are observed in the cross-subject and cross-session experiments respectively. Fine-granularity alignment encourages wider inter-class margin. As shown in Table VIII, the performance of domain alignment also depends on feature normalization, where the removal of the batch normalization layer leads to decreased performance. To improve the quality of pseudo class labels that the fine-grained alignment relies on, these pseudo labels are pre-filtered by a pre-determined threhod to remove the unreliable ones. The fine-grained adaptation is dynamically weighted using a dynamic coefficient, which is proportionate to the number of training iterations. By improving the domain discriminability of our model as the training proceeds, it is observed that more reliable pseudo labels for fine-grained analysis can be obtained.
Figure 5: Comparison of emotion recognition accuracies of different domain adaptation algorithms on SEED and SEED-IV datasets respectively for (a) Cross-subject experiment; and (b)Cross-session experiment. Our MS-DCDA model outperforms in all experiments three single-source methods: DCORAL [58], DAN [35], DANN [36], DDC [57], as well as two multi-source methods: MS-MDA [49], and MS-ADA [43].
@@
To examine impact of different brain lobes on emotion recognition, the 62 EEG electrodes are divided into four regions based on their location: front, parietal, temporal, and occidental. As shown in Table IX, the frontal and parietal lobes appear to be the most active brain lobes while watching emotional movie clips, followed by the occipital and temporal lobes. Although only six electrods are allocated for the temporal lobe, its high activity suggests that the expression of emotions is likely independent of the number of recognized electrode signals. The combination of frontal and parietal lobes consistently outperformed individual lobes for emotion recognition on both the SEED and SEED-IV datasets, reaffirming the important role of frontal and parietal lobes in emotional expression. In addition, as the temporal or occipital signals continue to increase and exceed a certain threshold, the recognition accuracy stops improving, and some may even appear to decrease. One possible explanation is that EEG signals measured from the occipital and temporal lobes exhibit relatively lower sensitivity to emotional states compared to those from the frontal and parietal lobes. Including the less sensitive occipital and temporal lobes may reduce the emotion recognition results. Despite of the adverse effect of the occupital and temporal lobe, the recognition accuracies remain higher in cases where brain lobes are combined compared to the recognition result from a single brain lobe. Our studies indicate that brain regions exhibit different sensitivities to emotional states, reaffirming the need to consider the combination of different lobes during emotion recognition.
VII Conclusion
In summary, a multi-source dynamic contrastive domain adaptation (MS-DCDA) method was introduced for EEG based emotion recognition. The proposed combines coarse- and fine-grained domain alignments, and dynamically optimize the domain transferability and discriminability, leading to improved model generalization, classification accuracies, wider inter-class margin, and higher intra-class compactness. The proposed MS-DCDA outperformed both the classical and the state-of-the-art methods on both the SEED and SEED-IV datasets by a large margin. In addition, Our study also suggests greater emotional sensitivity in the frontal and parietal brain lobes, providing insights for mental health interventions, personalized medicine, and development of preventive strategies. The broader significance of this research also lies in its potential applications across various fields. For instance, enhanced emotion recognition capabilities can improve human-computer interaction by allowing systems to respond more precisely to the emotional states of users.

### v2

@@ 5.1 Cross-Subject Results
The comparisons of the emotion recognition results of our MS-DCDA model and the baseline algorithms are presented in Tables 1, where the best results are highlighted in bold font. The alternative methods are tuned to their best performances. If the relevant parameters are published, we will use them. If not, the parameters will remain consistent with our method. The cross-subject experimental results on SEED and SEED-IV, which are based on both the state-of-the-art methods and our proposed method, are presented in Table 2. On the SEED dataset, the constructed MS-DCDA model significantly improves recognition performance, achieving 90.84% for the third session and 88.75% for the average of three sessions. In comparison to other methods, such as MS-MDA and MS-ADA, both being multiple source domain adaptation methods, our method demonstrates superior performance. Compared to MS-MDA which achieves the highest average accuracy, our method has improved accuracy by 3.01%. Compared to MS-ADA which achieves the highest average accuracy, our method has improved accuracy by 3.77%. On the SEED-IV dataset, our method achieved the highest accuracy 78.49% for the second session, which is 2.38% higher than MS-MDA, 74.34% for the average of three sessions, which is 3.12% higher than MS-MDA. Additionally, both F1 and Kappa scores of our model outperform the alternative methods.
@@
To further validate the superiority of our method, we compared it with another set of state-of-the-art methods. As illustrated in Table 3, our method outperforms other methods in mean accuracy for both the SEED and SEED-IV datasets. The multi-source domain adaptation algorithm MFA-LR, HVF2N-DBR, and MWACN also work well in the cross-subject experiments. Our algorithm dynamically balance the model domain transferability and discriminability and achieved higher accuracy than MFA-LR (uses class-level alignment loss) and MWACN (uses association reinforcement to adapt conditional distribution). The recognition accuracy is respectively increased by 1.51% and 3.50% in cross-subject experiments compared to the alternative methods (HVF2N-DBR for SEED dataset and MFA-LR for SEED-IV dataset).
5.2 Cross-Session Results
The specific subject results of the best session in cross-session experiments are shown in Tables 4. Our algorithm, which integrates dynamic multi-source domain adaptation and class alignment techniques, achieved the highest accuracy across most subjects. On the SEED dataset, some subjects achieved the highest accuracy of 100% in cross-session experiments. On the SEED-IV dataset, the subject achieved the highest accuracy of 95.55% and 94.98% in cross-session experiments. Notably, the results on the SEED dataset are significantly better than the results on the SEED-IV dataset. On the one hand, this may be because the SEED-IV dataset has more sentiment classification than the SEED dataset, which increases the difficulty of class feature recognition. On the other hand, it may be because the sample size of SEED-IV is much smaller than that of SEED. The cross-session experimental results of SEED and SEED-IV are provide in Table 5. Our MS-DCDA model outperformed the rest models for all four metrics.
5.3 Ablation Studies
Ablation on Loss Terms In this section, we conduct ablation experiments to investigate the different components and equilibrium factors of the model on performance. For the sake of conciseness and consistency, all ablation experiments were conducted on the SEED. The results of the ablation experiment for the model components in cross-subject and cross-session experiments are shown in Table 6. Compared to the accuracy of 88.75% for full model components in cross-subject experiments, the removal of the MMD loss leads to a decrease in accuracy by 0.74%, omitting the disc loss decreases accuracy by 2.1%, and excluding the SCD loss results in a decrease of 4.49%. It can be seen that SCD has the greatest performance improvement, while MMD has the smallest performance improvement. Even without using SCD, the accuracy remains higher than the baseline model by 0.13%. Compared to the accuracy of 93.92% for full model components in cross-session experiments, excluding the MMD loss results in a decreased accuracy of 0.51%, removing the disc loss reduces accuracy by 1.76%, and omitting the SCD loss leads to a decrease of 2.97%. Similar to the cross-subject experiments, SCD shows the most substantial performance improvement, followed by disc, and MMD. The accuracy of not using SCD is still higher than that of the baseline model by 0.27%. These results emphasize the indispensability of each loss component.
Ablation on Learning Strategies Table 7 presents the comparison of using dynamic weighting factor $\tau$ in Eq. 14 against several combinations of static weights. It can be seen that using dynamic weighting factor shows better performance. Using dynamic weighting factor $\tau$ promoting dynamic balance between domain alignment and class discrimination effectively avoids performance degradation caused by excessive pursuit of alignment or discriminability.
Table 6: Ablation Study on Loss Functions
@@ 5.3 Ablation Studies
        CE only       84.13$\pm$7.89       90.68$\pm$6.90
        w/o. MMD       88.01$\pm$9.95       93.41$\pm$5.87
        w/o. DISC       86.65$\pm$9.61       92.16$\pm$5.63
        w/o. SCD       84.26$\pm$7.66       90.95$\pm$5.00
        ALL       88.75$\pm$8.88       93.92$\pm$8.01
Table 7: Ablation Study on Static vs. Dynamic Learning
@@
Ablation on Layer Normalization The proposed MS- DCDA preserves the coarse-grained inter-domain alignment while improving the intra-domain and intra-class alignment through the fine-grained class-level adaptation. With the use of the SCD penalty, the increases in recognition accuracy of 4.49% and 2.97% are observed in the cross-subject and cross-session experiments respectively. Fine-granularity alignment encourages wider inter-class margin. As shown in Table 8, the performance of domain alignment also depends on feature normalization, where the removal of the batch normalization layer leads to decreased performance. To improve the quality of pseudo class labels the fine-grained alignment relies on, unreliable labels are pre-filtered by an empirical threshold. The fine-grained adaptation is dynamically weighted using a dynamic coefficient, which is proportionate to the number of training iterations. By improving the domain discriminability of our model as the training proceeds, it is observed that more reliable pseudo labels for fine-grained analysis can be obtained.
Table 8: Ablation Study on Layer Normalization (LN)
Method Cross-subject Cross-session
SEED SEED-IV SEED SEED-IV
CFE+LN 85.51$\pm$7.99 70.71$\pm$9.45 93.59$\pm$7.52 76.12$\pm$14.16
MBC+LN 82.86$\pm$7.49 70.89$\pm$11.07 90.06$\pm$7.21 77.46$\pm$12.32
BOTH+LN 88.75$\pm$8.89 74.34$\pm$10.62 93.92$\pm$8.01 77.53$\pm$12.13
Table 9: Ablation Study on Brain Lobes
Lobes Cross-subject Cross-session
SEED SEED-IV SEED SEED-IV
F 86.56$\pm$9.63 73.5$\pm$14.09 92.97$\pm$7.64 75.57$\pm$12.50
O 85.14$\pm$8.76 66.03$\pm$9.73 91.27$\pm$7.93 73.64$\pm$12.53
P 85.62$\pm$9.17 70.41$\pm$11.60 92.35$\pm$7.37 76.94$\pm$12.62
T 77.93$\pm$13.19 65.45$\pm$11.81 87.78$\pm$9.37 71.31$\pm$13.59
F+P 87.53$\pm$9.57 74.12$\pm$11.20 95.22$\pm$5.71 76.94$\pm$12.86
O+T 85.55$\pm$10.84 72.10$\pm$13.22 91.91$\pm$8.02 74.86$\pm$12.07
F+P+O 86.35$\pm$8.85 72.25$\pm$12.49 92.11$\pm$7.50 77.64$\pm$13.10
F+P+T 88.33$\pm$8.78 72.55$\pm$13.26 94.57$\pm$6.10 78.15$\pm$12.02
All 88.75$\pm$8.89 74.34$\pm$10.62 93.92$\pm$8.01 77.53$\pm$12.13
  • F: Frontal; O: Occipital; P: Parietal; T: Temporal
@@
Ablation on Brain Lobes To examine impact of different brain lobes on emotion recognition, the 62 EEG electrodes are divided into four regions based on their location: front, parietal, temporal, and occidental. As shown in Table 9, the frontal and parietal lobes appear to be the most active brain lobes while watching emotional movie clips, followed by the occipital and temporal lobes. Although only six electrodes are allocated for the temporal lobe, its high activity suggests that the expression of emotions is likely independent of the number of recognized electrode signals. The combination of frontal and parietal lobes consistently outperformed individual lobes for emotion recognition on both the SEED and SEED-IV datasets, reaffirming the important role of frontal and parietal lobes in emotional expression. In addition, as the temporal or occipital signals continue to increase and exceed a certain threshold, the recognition accuracy stops improving, and some may even appear to decrease. One possible explanation is that EEG signals measured from the occipital and temporal lobes exhibit relatively lower sensitivity to emotional states compared to those from the frontal and parietal lobes. Including the less sensitive occipital and temporal lobes may reduce the emotion recognition results. Despite of the adverse effect of the occipital and temporal lobe, the recognition accuracies remain higher in combined brain lobes compared to a single brain lobe. Our studies indicate that brain regions exhibit different sensitivities to emotional states, reaffirming the need to consider the combination of different lobes during emotion recognition.
5.4 Visual Analysis
Confusion Matrix Analysis. The confusion matrices of cross-subject and cross-session experiments for three sessions on SEED and SEED-IV are provide in Fig. 3. As can be seen in Fig. 3(a) and 3(b), our MS-DCDA method exhibits strongest sensitive to the positive emotion on the SEED data, with slightly reduced sensitivity in neutral and negative emtions. It is observed in Fig. 3(c) and 3(d) that neutral emotion is less likely being confused than fear and sadness on SEED-IV by our model, which may be explain by the fact that both fear and sadness are particularly negative and exhibit similarities in the activated electrode signals.
Figure 3: Confusion matrices analysis of our MS-DCDA model for (a) Cross-subject experiment on SEED dataset; (b) Cross-subject experiment on SEED-IV dataset; (c) Cross-session experiment on SEED; and (d) Cross-session experiment on SEED-IV. Sensitivity are the strongest to the positive and neutral emotions in SEED and SEED-IV dataset, with slightly reduced sensitivity to the rest emotions.
T-SNE Analysis. The T-SNE plots [66] of feature distributions are provide in Fig. 4 for SEED dataset, which is richer in source diversity. We randomly picked 100 EEG samples from each subject for visualization to display changes in feature distribution caused by model training. As shown in Fig. 4(a), the distribution of most of the samples of original features are observed to be concentrated in a singular area, with a few outliers among individual subjects. This distribution affirms that the distribution EEG features may overlap in low-level feature space. This observation becomes more clear with normalization, where a more uniform and centralized distribution appears. The feature distributions produced by the DAN model and the MS-DCDA model are shown in Fig. 4 (b) and Fig. 4 (c) respectively. Overlap between the source and target features are observed in the DAN results. On the other hand, our MS-DCDA represents each source domain and target domain feature distribution separately, showcasing similarity and overlap between the targets and each individual source domain. Fig. 4(d) shows the initial distribution of combined source domain features and target domain features. We also show the feature distribution of source ensemble and target learned by DAN and our MS-DCDA respectively. While maintaining the consistency in source domain, it can be observed in Fig. 4(e) that DAN does not show a clear trend of class cluttering. In contrast, our MS-DCDA model adapts T to individual sources and a complementary multi source ensemble (SE) with class-awareness. It also dynamically adjusts the weights of domain transferability and discriminability, leading to improved classification accuracies, wider inter-class margin, and higher intra-class compactness (see Fig. 4(f)).
Data Transfer Studies. The SEED and SEED-IV datasets are similar in experimental design, collection methods, data pre-processing, and analogical categories. They differ only in the number of emotion classes. Here we train our model with one dataset and test using the other as dataset transfer studies. To have the same number of emotion classes in each set, we merge the sad and fear classes in SEED-IV into a single negative class in SEED, and correspond neutral and happy in SEED-IV with the neutral and positive in SEED respectively. The results of the data transfer studies are shown in Fig. 5. It is observed that our MS-DCDA model outperforms alternative SS-DA and MS-DA algorithms in both cross-subject (see Fig. 5(a)) and cross-session experiments (see Fig. 5(b)), demonstrating an improved generalization ability. It is also found that the effect of data order on the performance of data transfer is minimal, especially in cross-session experiments, which may be explained by the similarity of the two datasets. The inter-class and intra-class alignment in our MS-DCDA model play important roles in extraction of fine-grained emotional features and enhancing the robustness of our model in various transfer scenarios.
6 Discussion
The presented MS-DCDA framework demonstrated improved emotion recognition performance on both SEED and SEED-IV datasets. In cross-subject experiments, we achieved an accuracy of 90.48% and 90.85% for S1 and S3 on the SEED dataset, with a relatively lower accuracy of 84.98% in S2. On the SEED-IV dataset, our MS-DCDA performs better in S2 with a leading 78.49% accuracy, while the accuracies are slightly reduced to 72.36% and 72.16% for S1 and S3 due to more complicated data dependencies in that session. The cross-session variations can result in variations in recognition accuracies if data are obtained through repetitive experiments. The robustness of the proposed framework against such variations was demonstrated through extensive cross-validation. The observation that cross-subject results appear to be less robust compared to the corresponding cross-session results may be explained by the individual subject disparities in EEG being more complicated than its non-stationary problems. As demonstrated in Tables 2, 3 and 5, multisource domain adaptation generally achieves superior performance than single-source methods for EEG-based emotion recognition tasks. The ablation studies on various loss functions are shown in Table 6, where the accuracy of our model improved from 84.13% and 90.68% by incorporating multi-source loss penalties. The use of pair-wise neural network branches allows adaptation of target to different source distributions through distinctive feature representations and improves the accuracy and generalization of our model. However, computational complexity may increase linearly with the number of sources. Dynamic routing among branches using a mixture-of-expert architecture may improve the recognition accuracy at a reduced training cost.
Figure 4: T-SNE illustration of domain adaptation and emotion recognition on SEED dataset. (a) Distribution of 14 individual source subjects (S1-S14), the ensemble of the 14 individual sources (S15, source ensemble) and a target subject. (b) Distribution of the single source (S) and the target learned by DAN [41]. (c) Distribution of the 14 individual source subjects and the target subject learned by our MS-DCDA. (d) The Distribution of the source ensemble and the target. (e) Distribution of the single source and the target learned by DAN. (f) Distribution of the source ensemble and the target learned by our MS-DCDA.